OpenStudy (anonymous):

Calculus Help!! If f is continuous and ∫(upper limit:9,lower limit:0) f(x)dx=8, find ∫(upper limit:3,lower limit:0) xf(x^2)dx.

OpenStudy (anonymous):

what should I do when they change the upper limit?

OpenStudy (zarkon):

\[u=x^2\]

OpenStudy (dumbcow):

u = x^2 du = 2x dx --> dx = du/2x integral 1/2 f(u) du, where limits for x are 0 to 3 but change limits for u from 0 to 9

OpenStudy (anonymous):

I still don't really get how to change the limit!!Can you show some steps,please?

OpenStudy (dumbcow):

original limits 0 < x < 3 u = x^2 0^2 < u < 3^2 0 < u< 9

OpenStudy (anonymous):

I got it!!Thanks a lot!!